• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Boeglin

    Log entry time 16:05:55 on May28,2008

    Entry number 240105

    keyword=Shift Summary

    Continue with run 1662

    9:40 MCC called they are looking to take Beam away at 10am to study their injector problems.
    11:07 looks like the beam is back start next run : 1665, still many trips (MCC has not yet called that they are finished with their studies)

    14:50 move to carbon target for optics measurement
    15:30 finished optics move to next kinematics
    Showercounter HV tripped reset

    Completed Runs

    Run Nr. Time Nr. of Events Comment
    1662 8:33 5 M
    1663 9:14 5 M
    1664 10:10 3.1 M MCC optimizing (last 24 min, no beam)
    1665 11:50 5 M
    1666 12:50 5 M
    1667 13:35 5 M
    1668 14:12 5 M
    1669 14:51 5 M completes kin 6 0%
    1670 15:23 0.4 M Pointing w/o raster
    1671 15:30 0.36 M Pointing with raster
    new kin: +2%
    1672 15:59 0.77 M test run for Shower counters
    ready to take data on new setting